a uniform metal rod is used as a bar pendulum if the room tempature rises by 100 and the coefficient of linear expansion of the metal of the rod is 2x10-6  the period of the pendulum will have percentage increase of?
1]2x10-3
2]3x10-3
3]2x103
4]1x10-3

Grade:8

1 Answers

Vikas TU
6 years ago
Fractional change in period =>
ΔTT=12αΔθ=12×2×10−6×10=10−5.
the period of the pendulum will have percentage increase of=>
% change=ΔTT×100=10−5×100=10−3